OpenStudy (anonymous):

A first order linear equation in the form y'+p(x)y=f(x) can be solved by finding an integrating factor mu (x) = exp( int (p(x) dx)) (1) Given the equation y' + 3 y = 2 find mu(x) = (2) Then find an explicit general solution with arbitrary constant C. y = . (3) Then solve the initial value problem with y(0)=2 y = .

myininaya (myininaya):

what is considered mu?

OpenStudy (ipwnbunnies):

Your integrating factor should be: \[\mu (x) = e^{\int\limits_{}^{} p(x) dx}\] This form of differential equation is written as: y' + p(x)*y = g(x)

myininaya (myininaya):

oh integrating factor is mu ?

OpenStudy (anonymous):

\[\mu (x)\] the integrating factor. I don't know how to start.

myininaya (myininaya):

@ipwnbunnies gave you the formula